what is answer for x+2+x=22
what is the answer for 9t-6-6t=6
what is the answer for -23= -2a-10+a
what is the answer for 5(2x-3)=15
what is the answer for 2(8+4c)=32

Answer by ReadingBoosters(3246)

Starting with x+2+x=22
2x + 2 = 22
2x = 20
x = 10
Starting with 9t-6-6t=6
3t - 6 = 6
3t = 12
t = 4
Starting with -23=-2a-10+a
-23 = -a - 10
-13 = -a
13 = a
Starting with 5(2x-3)=15
10x-15=15
10x=30
x=3
Starting with 2(8+4c)=32
16 + 8c = 32
8c = 16
c=2